Find the length of a rectangular lot with a perimeter of 84 m if the length is 6 m more than the width.


Answer by checkley77(12844)

L=W+6
2L+2W=84
2(W+6)+2W=84
2W+12+2W=84
4W=84-12
4W=72
W=72/4
W=18 ANS. FOR THE WIDTH.
L=18+6=24 ANS. FOR THE LENGTH.
PROOF:
2*24+2*18=84
48+36=84
84=84